OpenStudy (chacha1234):

Please help will fan and medal! Audra counted change collected for a class fundraiser. Audra wrote the equation 0.10x+0.25y=5.30 to represent the situation. What did she define as the variable x? value of a dime value of a quarter number of dimes collected number of quarters collected

OpenStudy (chacha1234):

@RhondaSommer @xGuardians

OpenStudy (chacha1234):

@Anaise

OpenStudy (chacha1234):

@AloneS

OpenStudy (chacha1234):

I'm so confused

OpenStudy (chacha1234):

I don't get it?

RhondaSommer (rhondasommer):

they gave you a calculator, which is useless in this equation

OpenStudy (narutoboy14):

for real

OpenStudy (chacha1234):

@RhondaSommer So can you help?

OpenStudy (chacha1234):

It's 2:00 am I'm super sleepy but I have to finish the couple of questions that are left.

RhondaSommer (rhondasommer):

0.10x+0.25 \[quarters~are~~\frac{25}{100}=0.25\]\[dimes~are~~~~~~\frac{10} {100}=0.10\] so if x is multiplied by (0.10) what is (x) going to be?

OpenStudy (anonymous):

0.10 dollar=10 cents=value of one dime so x is number of dimes.

OpenStudy (chacha1234):

dimes

RhondaSommer (rhondasommer):

perfect :D

OpenStudy (chacha1234):

Value of a dime*

OpenStudy (chacha1234):

One more please?

RhondaSommer (rhondasommer):

sure :D

OpenStudy (chacha1234):

Wait, is it number of dimes collected or value of a dime?

OpenStudy (chacha1234):

I think it's number of dimes collected..?

RhondaSommer (rhondasommer):

number of dimes. because it is being multiplied by the value of a dime

OpenStudy (chacha1234):

Ok

OpenStudy (chacha1234):

Carissa works as a babysitter during her summer vacation. She gets paid one rate for daytime hours and a higher rate for nighttime hours. One week, she worked 12 daytime hours and 4 nighttime hours and earned $120. The next week, she earned $60 for a total of 4 daytime hours and 3 nighttime hours. Let x represent the hourly daytime rate and y represent her hourly nighttime rate. 12x+4y=120 4x+3y=60 What is the solution to the system of equations? (3,4) (4, 12) (6, 12) (10, 30)

OpenStudy (chacha1234):

hey @xGuardians

OpenStudy (chacha1234):

Care to help?

OpenStudy (chacha1234):

@RhondaSommer Hello?

RhondaSommer (rhondasommer):

OH MY GOODNESS. BE FRICKIN PATIENT

OpenStudy (chacha1234):

IM SO SORRY!

OpenStudy (anonymous):

multiply second eq. by 3 and subtract from eq 1.

RhondaSommer (rhondasommer):

ok \[No. 1~~~~12x+4y=120\]\[No. 2~~~~4x+3y=60\]\[multiply~No.2~equation~by~-3\]\[(4x+3y=60)-3\]\[wich~becomes~-12x-9y=-180\]\[combine No.1~and~modified~No.2~divide~and~both~sides~by~-5\]\[~~~12x+4y=~~~120\]\[-12x-9x=-180\]\[---------\]\[~~~~~~~~~~\frac{-5y}{-5}=~\frac{-60}{-5}\]\[~~~~~~~~~~~~~~~~~y=~~12\]\[place~12~in~for~(y)~in~the~origonal~No.2~equation\]\[4x+3(12)=60\]\[4x+36=60\]\[subtract~36~from~both~sides~and~divide~both~!sides~by~4\]\[\frac{4x}{4}~~~~~~~=\frac{24}{4}\]\[x~~~~~~~~~~~=6\]

OpenStudy (chacha1234):

(6,12)?

OpenStudy (chacha1234):

Yes!

RhondaSommer (rhondasommer):

so if \[(x,y)\]and\[x=6~y=12\]\[(6,12)\]

OpenStudy (chacha1234):

Thanks so much!

OpenStudy (chacha1234):

How can I repay you?

RhondaSommer (rhondasommer):

yeah :D sorry it took a while to write all that in. :D

OpenStudy (chacha1234):

I already fan and medaled anything else?

RhondaSommer (rhondasommer):

uh...that's it...maybe right a testimonial for me?

OpenStudy (chacha1234):

Already did I think..?

RhondaSommer (rhondasommer):

if you hover over my name you can right a testimonial for them...

OpenStudy (chacha1234):

I did

OpenStudy (chacha1234):

Let me see my grade

OpenStudy (chacha1234):

90% thanks so much

RhondaSommer (rhondasommer):

:D yeah no problem :D